### v1

# Universal scaling law for quantum droplet formation

Abstract
Given the right set of circumstances, ultracold quantum gases are able to change character and condense into a liquid state of quantum droplets. The size distribution of the droplets is determined dynamically in the condensation process. A semi-quantitative argument is presented which suggests that, at zero temperature, a multiple droplet system has is a preferred scale $\propto v^{-d}$, with $d\approx 0.5$, where $v$ is the rate of change of parameters at the time of droplet formation. Numerical simulations of two dimensional systems strongly support a power law, but with an exponent $d\in(0.3,0.4)$.
Introduction: Examples of quantum fluids such as liquid helium and ultracold quantum gases are characterised as either liquids or gases. Remarkably, given the right set of circumstances, ultracold quantum gases are able to change character and condense into a liquid state of quantum droplets Petrov (2015); Petrov and Astrakharchik (2016). This new physical phenomenon opens up an avenue into studying quantum liquids in a weakly interacting system Semeghini et al. (2018); Cheiney et al. (2018); Cabrera et al. (2018); Ferioli et al. (2019). Furthermore, it has been suggested Moss (2024) that droplets are non-relativistic analogues of relativistic “oscillon” phenomena of interest in elementary particle physics Makhankov et al. (1979); Gleiser et al. (1991); Gleiser (1994).
The existence of liquid droplets requires a balance between the attractive forces in an unstable atomic gas, and some kind of stabilising force. One possible stabilising force arises from quantum vacuum fluctuations in an atomic mixture Petrov (2015); Petrov and Astrakharchik (2016). The quantum forces are small, but can play an important role when other mean field forces almost cancel out. Examples of droplet formation have been seen in Bose-Einstein condensates with mixtures of two states with intercomponent attraction and intracomponent repulsion, where the vacuum energy contribution is repulsive and stabilising Semeghini et al. (2018); Cheiney et al. (2018); Cabrera et al. (2018); Ferioli et al. (2019).
This paper aims to answer a simple question about the dynamics of multiple droplet formation, what governs the size and separation of the droplets? It will be argued that, changing the parameters of a suitable droplet-forming system through the droplet forming regime at a rate $v$ leads to droplets of characteristic size $r_{d}\propto v^{-d}$. This result is valid up to a limiting rate, and depends on there being a large number of droplets. It is conjectured that the scaling law is independent of the detailed form of the quantum corrections, and consequently independent of dimension. A semi-quantitative argument for the scaling law is presented, and this is backed up by numerical simulations in two dimensions. So far, experiments have only been done with one or two droplets, where droplet size is determined by the total number of atoms, and so new experiments would be needed to test the new theoretical prediction.
The scaling law for droplet sizes resembles the Kibble-Zurek scaling law for structure formation in second order phase transitions Kibble (1976); Zurek (1985); del Campo and Zurek (2014), which involves a power of the rate parameter. There is also good evidence that a Kibble-Zurek type of scaling law applies to some first order phase transitions in cold atom systems Panagopoulos and Vicari (2015); Kang et al. (2017); Liang and Zhong (2017); Coulamy et al. (2017); Qiu et al. (2020). The physical origin of the scaling laws for the droplets has some similarities, in that there is a limitation on the size of structures that can form on some timescale. In the thermal case, this is set by the relaxation time of the medium. In the case of droplet formation, which can occur at zero temperature in the absence of dissipation, this is related to the growth rate of the droplet instability.
Droplet formation: Consider a condensate of two atomic states with atom number densities $n_{\uparrow}$ and $n_{\downarrow}$, and scattering lengths $a_{\uparrow\uparrow}$, $a_{\downarrow\downarrow}$ and $a_{\uparrow\downarrow}$. The mixture becomes unstable when $a_{\uparrow\uparrow}a_{\downarrow\downarrow}-a_{\uparrow\downarrow}^{2}<0$. The case of “balanced” number densities, where $n_{\uparrow}/n_{\downarrow}=(a_{\downarrow\downarrow}/a_{\uparrow\uparrow})^{1/2}$, is favoured on energy grounds. The theory in this case can be described by a single mean field $\psi$ with total atom density $n$ and quantum-corrected energy density $V_{LHY}(n)$ Lee et al. (1957). The field $\psi$ satisfies the Gross-Pitaevsky equation Petrov (2015); Petrov and Astrakharchik (2016),
$i\hbar\dot{\psi}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi-\mu\psi+V_{LHY}^{\prime}(n)\psi.$ (1)
Stability of a homogeneous mixture with density $\bar{n}$ requires $V_{LHY}^{\prime\prime}(\bar{n})>0$, and the onset of instability is marked by a density parameter $n_{d}$ such that $V_{LHY}^{\prime\prime}(n_{d})=0$  [^1] The droplet density depends on atomic scattering lengths. Instability can be approached either by increasing $n_{d}$ or decreasing the mean density $\bar{n}$.. The final density inside a droplet is equal to $n_{d}$ up to a factor of Euler’s constant.
The chemical potential parameter is a simple convenience that allows us to fix $\dot{\psi}=0$ for the homogeneous system, setting $\mu=V_{LHY}^{\prime}(\bar{n})$. The stability of the homogeneous state is analysed in perturbation theory by introducing the Bogoliubov modes, which have dispersion relation
$\omega^{2}=\frac{k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2\bar{n}V_{LHY}^{\prime\prime}\right).$ (2)
Thus, unstable modes exist when $n_{d}>\bar{n}$ and $V^{\prime\prime}_{LHY}$ becomes negative. Since these instabilities mark the onset of droplet formation, it seems reasonable to suppose their wavelengths will be related to the distance between droplets. The growth rate of these modes, $v_{d}={\rm Im}(\omega)$ is a function of wave number and time. It has a maximum at a given time for a particular wave number $k_{\rm max}$,
$v_{d}=\frac{\hbar k_{\rm max}^{2}}{2m}=\frac{1}{\hbar}|\bar{n}V_{LHY}^{\prime\prime}|.$ (3)
Now consider what happens to a uniform system that starts in the stable regime with $n_{d}<\bar{n}$, following which the physical parameters are slowly changed so that $n_{d}$ increases at a rate $v$. Initially, the growth rate $v_{d}=0$, and there is no droplet formation. When $n_{d}\geq\bar{n}$, the growth rate $v_{d}$ slowly increases from zero. The growth rate becomes comparable to the system timescale when $v_{d}\sim v$. Taking $v_{d}=v$ sets a critical length scale $2\pi/k_{\rm max}$ from Eq. (3), which we identify as the preferred separation between neighbouring droplets $r_{\rm sep}$. The conclusion is that the droplet separations scale with the rate $v$ as
$r_{\rm sep}\propto v^{-1/2}.$ (4)
Knowing the mean density of the system $\bar{n}$, the density of a typical droplet $n_{d}$ and the separation between droplets fixes the droplet radius $r_{d}$, $r_{d}=r_{\rm sep}\sqrt{\bar{n}/\pi n_{d}}\propto v^{-d}$. There is a caveat for large rates $v$, when the right hand term in Eq. (3) can have a maximum value $v_{\rm max}<v$, in which case the droplet separations would be set by $v_{\rm max}$ and would become independent of $v$.
The scaling rule governs droplet radii once they form. After formation, droplet mergers can have a significant effect on their sizes, particularly when their separation is small. Furthermore, three particle collision losses can cause droplets to evaporate and reduce in size Kim and Zubarev (2004). Droplet mergers are included in the numerical results discussed below.
Simulation: A two-dimensional numerical simulation has been used to test the theoretical prediction for the droplet radii. Near the unstable parameter range, the Gross-Pitaevski equation (GPE) becomes Petrov and Astrakharchik (2016)
$i\hbar\dot{\psi}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi-\mu\psi+g^{\prime}n\psi+gn\left(\ln\frac{n}{n_{0}}-1\right)\psi.$ (5)
where $n_{0}$ is a fixed parameter. Classical and quantum terms have been separated out and given coefficients $g^{\prime}$ and $g$ respectively. The formation of quantum droplets happens when an attractive classical force term ($g^{\prime}<0$) balances a repulsive quantum term. In practice, such a two dimensional system can be embedded into three dimensions with transverse dimension $a_{\perp}$, in which case the parameters $g^{\prime}$ and $g$ are related to the scattering lengths by
$g^{\prime}=\frac{\sqrt{8\pi}\hbar^{2}}{ma_{\perp}}\frac{a_{\uparrow\uparrow}a_{\downarrow\downarrow}-a_{\uparrow\downarrow}^{2}}{a_{\uparrow\uparrow}+a_{\downarrow\downarrow}-2a_{\uparrow\downarrow}},\qquad g=\frac{2\hbar^{2}}{m}\frac{|a_{\uparrow\downarrow}|^{2}}{a_{\perp}^{2}},$ (6)
and $n_{0}$ is of order $1/(4\pi a_{\perp}a_{\uparrow\downarrow})$ Ilg et al. (2018); Pelayo et al. (2025). Note that, after combining the potential terms as in Eq. (1), the droplet density parameter $n_{d}=n_{0}e^{-g^{\prime}/g}$ inherits a very strong dependence on the value of $a_{\uparrow\uparrow}a_{\downarrow\downarrow}-a_{\uparrow\downarrow}^{2}$.
If we define the healing length $\xi_{0}=\hbar/\sqrt{mgn_{0}}$ as a natural length scale, and $\omega_{0}=gn_{0}/\hbar$, then the quantitative features of the GPE are governed by a single parameter $g^{\prime}/g$. On the other hand, the initial conditions depend on $\bar{n}/n_{0}$ and the fluctuations in the field depend on a parameter $n_{0}\xi_{0}^{2}$.
The GPE was solved using a Truncated Wigner approach, where Gaussian random initial conditions represent a vacuum or thermal state. Initial conditions were imposed in the regime $n_{d}<\bar{n}$. The value of $g^{\prime}$ was then ramped down slowly to put the system into the range $n_{d}>\bar{n}$ when droplets are able to form. The rate $v$ is defined by
$v=\left|\frac{d\ln n_{d}}{dt}\right|_{n_{d}=\bar{n}}=\frac{1}{g}\frac{dg^{\prime}}{dt}.$ (7)
Each run was stopped at the same final atom density $n_{d}$. The droplet sizes and separations were then extracted by implementing a search algorithm to identify the individual droplets.
Figure 1: The pictures show the density in a numerical (vacuum) simulation for the parameters given later in table 1. The top picture uses a lower rate $v=0.004{\rm ms}^{-1}$, and the bottom a rate four times larger $v=0.016{\rm ms}^{-1}$. The droplet size is noticeably larger at the lower rate, though mergers are present in both cases. Red dots show the droplet positions found using a search algorithm.
Two sample runs are shown in figure 1 with two different values for the rate $v$. The figures clearly show that the droplets in the upper picture with a smaller rate have a larger size. There is also a clear indication that droplets sometimes merge into larger droplets, and take on non-spherical shapes as a result.
The relation between droplet sizes and parameter evolution rates is shown in figure 2. The values of the parameters $g^{\prime}/g$ and $n_{0}\xi_{0}^{2}$ differ by factors up to 4 between the plots to show that the scaling relation goes beyond a trivial rescaling invariance of the GPE. The data are excellent fits to power-law relations $r_{d}\propto v^{-d}$. The small variation in the value of $d$ for large changes in regime is consistent with small effects due to droplet mergers. However, the power $d\in(0.3,0.4)$ is significantly smaller than the naive prediction $d=0.5$ given earlier.
Figure 2: Droplet sizes from the simulations plotted as a function of parameter evolution rate $v$. Plots (a) and (b) use the parameters in table I, whilst plots (c) and (d) use half the trapping frequency, $\omega_{\perp}=7.5\times 2\pi{\rm kHz}$. In (a) and (c), the points use different initial scattering lengths with fixed ramp duration. In (b) and (d), the initial scattering lengths are fixed whilst reducing the ramp duration. The lines show least squares fits $r_{d}\propto v^{-d}$, with (a) $d=0.327$, (b) $d=0.345$, (c) $d=0.375$ and (d ) $d=0.343$. The error bars show standard deviations over 10 runs.
Experimental realisation: The numerical simulations were based on a hypothetical BEC mixture of potassium-39 in two hyperfine states $|1,0\rangle$ and $|1,-1\rangle$. At a field of around $56.85{\rm G}$ a Feshbach resonance affecting the scattering length $a_{\uparrow\uparrow}$ satisfies the instability condition $g^{\prime}=0$ D’Errico et al. (2007); Roy et al. (2013); Tanzi et al. (2018). The parameters used for the simulations are given in table 1. The initialisation was done with $B>56.85{\rm G}$ and $B$ was reduced to take the system through $g^{\prime}=0$ at a rate $v$.
An experimental realisation of the system would be subject to thermal effects. In the first place, the quantum potential has been calculated using the quantum vacuum energy. Thermal corrections would change the result if the thermal energy per particle exceeded the ground state energy of the Bogoliubov modes, which happens at a temperature $T_{LHY}$ given in table 1. A more exacting limit comes from the initial quantum fluctuations used in the Truncated Wigner approach. Thermal fluctuations become important at temperatures above $T_{GPE}$.
Table 1: Physical parameters used in the potassium-39 simulations (for figures 2(a) and 2(b)).
. Parameter Value initial density $\bar{n}=50\mu m^{-2}$ system size $232\mu{\rm m}$ ramp time $543{\rm ms}$ rate $v\in(0.004,0.02){\rm ms}^{-1}$ critical field strength $B=56.85{\rm G}$ field time derivative $dB/dt\in(0.3,0.12){\rm mG}\,{\rm ms}^{-1}$ trap frequency $\omega_{\perp}=15\times 2\pi{\rm kHz}$ droplet density scale $n_{0}=201\mu m^{-2}$ final droplet density $n_{d}=820\mu m^{-2}$ healing length $\xi_{0}=(\hbar/mgn_{0})^{1/2}=2.34\mu m$ frequency scale $\omega_{0}=gn_{0}/\hbar=295{\rm Hz}$ LHY temperature limit $T_{LHY}=|g_{12}|n_{d}/k_{B}=279{\rm nK}$ vacuum temperature limit $T_{GPE}=\hbar\omega_{0}/k_{B}=2.4{\rm nK}$
Conclusions: Ultra cold quantum gases condense into liquid quantum droplets under certain parameter regimes. In a many droplet system, the size distribution of these droplets is determined dynamically in the droplet transition. It has been argued here that there is a refered droplet size and separation fixed in terms of the rate of change of parameters $v$ by $r_{d}\propto v^{-d}$, where $d\approx 1/2$. This is independent of the detailed nature of the system.
Numerical simulations show strong evidence for a power law, but with $d\in(0.3,0.4)$. It would be interesting to pursue this further beyond two dimensions, to examine the effects of thermal fluctuations, and to consider “unbalanced” mixtures Flynn et al. (2023). The proposed range of experimental parameters look challenging, but further numerical analysis may allow some movement in these values
The droplet transtion has similarities to other physical processes. In particular, the droplets are non-relativistic analogues of the “oscillons” in elementary particle physics Makhankov et al. (1979); Gleiser et al. (1991); Gleiser (1994). Oscillons play a role in phase transitions, and are a candidate for dark matter Moss (2024). The study of quantum droplets may shed light on these related areas of physics.
Data supporting this publication are openly available under a Creative Commons CC-BY-4.0 Licence in Dat (2025)
Acknowledgements.
The author is grateful for helpful comments from Tom Billam, Matt Johnson, Markus Oberthaler and Yansheng Zhang. This work is supported by the UK Science and Technology Facilities Council [grants ST/T00584X/1 and ST/W006162/1]. The author is also grateful for the hospitality of Perimeter Institute where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
Appendix A Supplementary Material
Renormalisation It is instructive to analyse the renormalisation of the quantum vacuum energy in two dimensions which has been used in the main text. We start with separate number densities in the two BEC components $n_{\uparrow}$ and $n_{\downarrow}$. The quantum vacuum energy density $\Delta V$ depends on the two-dimensional couplings $g_{ij}$. It requires a momentum cutoff $\kappa$ to obtain a finite resultPetrov and Astrakharchik (2016),
$\Delta V=\frac{\hbar^{2}}{m}\kappa^{4}+\frac{m}{8\pi}\kappa^{2}\sum_{\pm}c_{\pm}^{2}+\frac{1}{8\pi}\frac{m^{3}}{\hbar^{2}}\sum_{\pm}c_{\pm}^{4}\ln\frac{m^{2}c_{\pm}^{2}}{\hbar^{2}\kappa^{2}}$ (8)
where
$2mc_{\pm}^{2}=g_{\uparrow\uparrow}n_{\uparrow}+g_{\downarrow\downarrow}n_{\downarrow}\pm[(g_{\uparrow\uparrow}n_{\uparrow}-g_{\downarrow\downarrow}n_{\downarrow})^{2}+4g_{\uparrow\downarrow}^{2}n_{\uparrow}n_{\downarrow}]^{1/2}$ (9)
The first term in $\Delta V$ is the usual energy density divergence, and the second can be absorbed into a homogeneous phase rotation. The cutoff dependence $\ln\kappa^{2}$ in the final term depends on the combination
±$\sum_{\pm}c_{\pm}^{4}=\frac{1}{2m^{2}}\left(g_{\uparrow\uparrow}^{2}n_{\uparrow}^{2}+g_{\downarrow\downarrow}^{2}n_{\downarrow}^{2}+2g_{\uparrow\downarrow}^{2}n_{\uparrow}n_{\downarrow}\right).$ (10)
This has the same form as the classical scattering terms, and can be absorbed by renormalisation of the couplings. We put a cutoff dependence into the “bare” couplings and define the renormalised couplings $g_{ij}^{R}$ by
$g_{ij}(\kappa)=g_{ij}^{R}+\frac{1}{8\pi}\frac{m}{\hbar^{2}}(g_{ij}^{R})^{2}\ln(\kappa^{2}/\mu_{R}^{2}).$ (11)
The renormalisation scale $\mu_{R}$ must be introduced for consistency, and it replaces the cutoff $\kappa$ in Eq. (8). In some respects, $\mu_{R}$ is an arbitrary scale. However, other quantum corrections will depend on the renormalisation point, and some choices are better than others at suppressing these. In this example, one obvious source of corrections is the third dimension, and we can use results that include a third, compact, dimension to guide the choice of renormalisation point Ilg et al. (2018); Pelayo et al. (2025). These results contain a term similar to (8), with $\mu_{R}\sim 1/a_{\perp}$, which suggests choosing this as the renormalisation point.
Returning to the balanced case for the number densities, then both $c_{-}^{2}$ and the classical potential are $\propto g_{\uparrow\uparrow}g_{\downarrow}-g_{\uparrow\downarrow}^{2}$, which is small close to the instability point. The quantum term is also reduced by extra factors of $1/a_{\perp}$, so we neglect $c_{-}^{4}$, but keep the classical terms and $c_{+}^{4}$. By choosing $\mu_{R}$ as indicated above, we get the simple form for the GPE given in the main text.
Details of the numerical simulation The numerical scheme integrates the projected GPE (PGPE) with a variable timesstep integration scheme. The projection is introduced for de-aliasing, removing modes with $k$ values larger than half the lattice cutoff. The runs used to produce the resuts have grid size $512\times 512$, but the results are unchanged when the grid is increased to $1024\times 1024$.
Following the truncated Wigner approach, multiple runs start from initial conditions drawn from a gaussian distribution which replicates the symmetrised two-point correlation function of the quantised field fluctuations. The runs start in a regime with $g^{\prime}\neq 0$, and the correlation function
$\langle\{\delta\bar{\psi}(k)\delta\psi(k^{\prime})\}\rangle=\frac{1}{2}\frac{K+g^{\prime}\bar{n}}{\sqrt{K(K+2g^{\prime}\bar{n})}}\delta_{kk^{\prime}},$ (12)
[…]
Figure 3: The histogram shows the probability distribution of atom numbers in each lump. (Compiled from five separate runs with $v=0.01{\rm ms}^{-1}$.)
The value of $g^{\prime}$ in each run varies linearly for the ramp duration and ends at a fixed value of $g^{\prime}/g$, so that the final droplet density $n_{d}$ is fixed, The droplets are identified using a search algorithm. Figure 3 shows the atom number distribution over the droplets for a sequence of runs with the same parameters. There is a distinct separation into liquid droplets, on the right, with constant internal density, and small solitonic style lumps with radii of order the healing length. These small lumps are removed before testing the scaling law by putting a threshold $n_{d}/2$ on the internal density. The mean radius $r_{d}$ is defined from the atom numbers and density, and the separation $r_{\rm sep}$ from the number of droplets per unit area. Both satisfy identical scaling laws.
Details of the experimental proposal The scattering length $a_{\uparrow\uparrow}$ of the hyperfine state $|1,0\rangle$ of potassium-39 varies strongly with magnetic field around $57{\rm G}$, and $a_{\uparrow\uparrow}a_{\downarrow\downarrow}-a_{\uparrow\downarrow}^{2}$ changes sign from positive (stable) a little above $56.85{\rm G}$ to negative below. The rate $v$ used in the text can be related to the rate of change of the magnetic field using Eq. (6) and scattering lengths from the literature D’Errico et al. (2007); Roy et al. (2013); Tanzi et al. (2018),
$v=\frac{1}{g}\frac{dg^{\prime}}{dt}\approx 131\frac{a_{\perp}}{1\mu m}\frac{dB}{dt}.$ (13)
Broadening the range of some of the parameters in table 1 is restricted by a number of constraints. Primary requirements are $r_{\rm sep}>r_{d}$, so that the droplets are well-separated. A second requirement $r_{d}>\xi_{0}$, so that the radius is larger than the thickness of the wall of the droplet. A third requirement is that the separation should be less than the overall size of the system. A last inequality $\xi_{0}>a_{\perp}$, is required for an effective two-dimensional theory. This last constraint may be relaxed when investigating systems which are not two dimensional.

### v2

@@
 The size distribution of the droplets is determined dynamically in the condensation process.
-A semi-quantitative argument is presented which suggests that, at zero temperature, a multiple droplet system has is a preferred scale $\propto v^{-d}$, with $d\approx 0.5$, where $v$ is the rate of change of parameters at the time of droplet formation.
-Numerical simulations of two dimensional systems strongly support a power law, but with an exponent $d\in(0.3,0.4)$.
+A semi-quantitative argument is presented which suggests that, at zero temperature, a multiple droplet system has is a preferred scale $\propto v^{-1/3}$, where $v$ is the rate of change of parameters at the time of droplet formation.
+Numerical simulations of two dimensional systems strongly support a power law $v^{-d}$, with an exponent $d\in(0.327,0.375)$.
 Introduction: Examples of quantum fluids such as liquid helium and ultracold quantum gases are characterised as either liquids or gases.
@@
 The scaling law for droplet sizes resembles the Kibble-Zurek scaling law for structure formation in second order phase transitions Kibble (1976); Zurek (1985); del Campo and Zurek (2014), which involves a power of the rate parameter.
-There is also good evidence that a Kibble-Zurek type of scaling law applies to some first order phase transitions in cold atom systems Panagopoulos and Vicari (2015); Kang et al.
+The physical origin of the scaling laws for the droplets has some similarities, in that there is a limitation on the size of structures that can form on some freeze-out timescale.
+In the thermal case, this is set by the relaxation time of the medium.
+In the case of droplet formation, which can occur in the absence of dissipation, the process is similar to the “underdamped” Kibble-Zurek effect Suzuki and Zurek (2024).
+However, an important difference is that droplet formation can happen at zero temperature from purely quantum fluctuations that cannot be modelled by a Langevin equation approach.
+There is also good evidence that a Kibble-Zurek type of scaling law applies to other situations, for example some first order phase transitions in cold atom systems Panagopoulos and Vicari (2015); Kang et al.
 (2017); Liang and Zhong (2017); Coulamy et al.
@@
 (2020).
-The physical origin of the scaling laws for the droplets has some similarities, in that there is a limitation on the size of structures that can form on some timescale.
-In the thermal case, this is set by the relaxation time of the medium.
-In the case of droplet formation, which can occur at zero temperature in the absence of dissipation, this is related to the growth rate of the droplet instability.
 Droplet formation: Consider a condensate of two atomic states with atom number densities $n_{\uparrow}$ and $n_{\downarrow}$, and scattering lengths $a_{\uparrow\uparrow}$, $a_{\downarrow\downarrow}$ and $a_{\uparrow\downarrow}$.
@@
 $\omega^{2}=\frac{k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2\bar{n}V_{LHY}^{\prime\prime}\right).$ (2)
-Thus, unstable modes exist when $n_{d}>\bar{n}$ and $V^{\prime\prime}_{LHY}$ becomes negative.
-Since these instabilities mark the onset of droplet formation, it seems reasonable to suppose their wavelengths will be related to the distance between droplets.
-The growth rate of these modes, $v_{d}={\rm Im}(\omega)$ is a function of wave number and time.
-It has a maximum at a given time for a particular wave number $k_{\rm max}$,
-$v_{d}=\frac{\hbar k_{\rm max}^{2}}{2m}=\frac{1}{\hbar}|\bar{n}V_{LHY}^{\prime\prime}|.$ (3)
-Now consider what happens to a uniform system that starts in the stable regime with $n_{d}<\bar{n}$, following which the physical parameters are slowly changed so that $n_{d}$ increases at a rate $v$.
-Initially, the growth rate $v_{d}=0$, and there is no droplet formation.
-When $n_{d}\geq\bar{n}$, the growth rate $v_{d}$ slowly increases from zero.
-The growth rate becomes comparable to the system timescale when $v_{d}\sim v$.
-Taking $v_{d}=v$ sets a critical length scale $2\pi/k_{\rm max}$ from Eq.
-(3), which we identify as the preferred separation between neighbouring droplets $r_{\rm sep}$.
-The conclusion is that the droplet separations scale with the rate $v$ as
-$r_{\rm sep}\propto v^{-1/2}.$ (4)
-Knowing the mean density of the system $\bar{n}$, the density of a typical droplet $n_{d}$ and the separation between droplets fixes the droplet radius $r_{d}$, $r_{d}=r_{\rm sep}\sqrt{\bar{n}/\pi n_{d}}\propto v^{-d}$.
-There is a caveat for large rates $v$, when the right hand term in Eq.
-(3) can have a maximum value $v_{\rm max}<v$, in which case the droplet separations would be set by $v_{\rm max}$ and would become independent of $v$.
+Thus, unstable modes exist when $n_{d}>\bar{n}$ and $V^{\prime\prime}_{LHY}$ becomes negative, marking the onset of droplet formation.
+If the parameters change at a rate $v$ near the droplet transition, then we can replace $V^{\prime\prime}_{LHY}$ by $-gvt$, where $g$ is related to the third derivative of $V_{LHY}$ and the instability sets in at $t=0$.
+Following Sazuki and Zurek Suzuki and Zurek (2024), finding the characteristic length scales can be broken down into two steps.
+The first step is to find the freeze-out time when the instability becomes non-linear, and the second step is to convert this into a length scale.
+Taking the dispersion relation (2) as a starting point, the Bogoliubov modes will satisfy a non-linear equation of the form
+$m\delta\ddot{\psi}-gnk^{2}vt\delta\psi+O(\delta\psi^{2})=0,$ (3)
+in the small $k$ limit.
+The $v$ dependence can be removed by introducing a time coordinate $\tilde{t}=v^{1/3}t$, and rescaling $\delta\psi$ to remove any $v$ dependence in the $\delta\psi^{2}$ terms.
+The solutions for any given $k$ mode are independent of $v$ and become non-linear at some fixed value for the rescaled time $\tilde{t}(k)$.
+Returning to the dispersion relation (2), there is a value of $k=k_{\rm max}$ for which the growth rate $\Im(\omega)$ reaches a maximum [^2] The freeze-out time was found using a long wavelength approximation which is relaxed when finding the length scale.
+This is consistent with the approach used in Ref.
+Suzuki and Zurek (2024), $k_{\rm max}\propto|V_{LHY}^{\prime\prime}|^{1/2}\propto(vt)^{1/2}$.
+We may interpret the corresponding length scale as the droplet separation $r_{s}$, $r_{s}\propto(vt)^{-1/2}$.
+Since the freeze-out time $t=v^{-1/3}\tilde{t}$, it follows that
+$r_{s}\propto v^{-1/3}.$ (4)
+Knowing the mean density of the system $\bar{n}$, the density of a typical droplet $n_{d}$ and the separation between droplets fixes the droplet radius $r_{d}$, $r_{d}=r_{\rm sep}\sqrt{\bar{n}/\pi n_{d}}\propto v^{-1/3}$.
 The scaling rule governs droplet radii once they form.
@@
 The small variation in the value of $d$ for large changes in regime is consistent with small effects due to droplet mergers.
-However, the power $d\in(0.3,0.4)$ is significantly smaller than the naive prediction $d=0.5$ given earlier.
+However, the power $d\in(0.327,0.375)$ is consistent with the naive prediction $d=1/3$ given earlier.
 Figure 2: Droplet sizes from the simulations plotted as a function of parameter evolution rate $v$.
@@
 In a many droplet system, the size distribution of these droplets is determined dynamically in the droplet transition.
-It has been argued here that there is a refered droplet size and separation fixed in terms of the rate of change of parameters $v$ by $r_{d}\propto v^{-d}$, where $d\approx 1/2$.
+It has been argued here that there is a refered droplet size and separation fixed in terms of the rate of change of parameters $v$ by $r_{d}\propto v^{-d}$, where $d=1/3$.
 This is independent of the detailed nature of the system.
-Numerical simulations show strong evidence for a power law, but with $d\in(0.3,0.4)$.
+Numerical simulations show strong evidence for a power law, with $d\in(0.327,0.375)$.
 It would be interesting to pursue this further beyond two dimensions, to examine the effects of thermal fluctuations, and to consider “unbalanced” mixtures Flynn et al.
@@
 The study of quantum droplets may shed light on these related areas of physics.
-Data supporting this publication are openly available under a Creative Commons CC-BY-4.0 Licence in Dat (2025)
+Data supporting this publication are openly available under a Creative Commons CC-BY-4.0 Licence in Dat
 Acknowledgements.
-The author is grateful for helpful comments from Tom Billam, Matt Johnson, Markus Oberthaler and Yansheng Zhang.
+The author is grateful for helpful comments from Tom Billam, Matt Johnson, Markus Oberthaler, Yansheng Zhang and Mark Hindmarsh.
 This work is supported by the UK Science and Technology Facilities Council [grants ST/T00584X/1 and ST/W006162/1].
-The author is also grateful for the hospitality of Perimeter Institute where part of this work was carried out.
+The author is also grateful for hospitality of the Perimeter Institute and the University of Helsinki where part of this work was carried out.
 Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
-Appendix A Supplementary Material
+Supplementary Material
 Renormalisation It is instructive to analyse the renormalisation of the quantum vacuum energy in two dimensions which has been used in the main text.